# Short-time simulation of quantum dynamics by Pauli measurements

### v1

@@ III Concrete example and reducing terms
To gain better intuition of how the expansion of time dynamics into a linear combination of Pauli operators of Eq. (5) might look like for a specific Hamiltonian, let us consider a Heisenberg Hamiltonian with equally weighted interactions, Take for example a one-dimensional Heisenberg Hamiltonian with equally weighted interactions, i.e.,
@@
Naively, we have upper-bounded the number of terms present in the expansion of the $K$-th order truncated Taylor series of the time evolution operator as $m_{\mathrm{tot}}=\sum_{k=0}^{K}L^{k}=(L^{K+1}-1)/(L-1)$. However, since the Pauli operators square to the identity and also form a group, many terms will collapse to the identity matrix or reproduce already existing terms as early as in the second order. Explicitly, for
@@
Thus, using algebraic properties of the Pauli operators, the total number of contributing terms for a Taylor series truncated after the second order is smaller than the worst-case bound of $(27(n-1)^{3}-1)/(3(n-1)-1)$ by $(n+9)(n-1)$, leading to $m_{\mathrm{tot}}=8n^{2}-23n+16$ terms. Note, however, that although these terms do not contribute to the number of Pauli operators, they still contribute their pre-factors which must be considered. These pre-factors $\gamma_{j}$ in Eq. (5) will then be sums of powers of $Jt$.
@@
$H=\sum_{i=1}^{n-1}\frac{J}{2}\mathrm{SWAP}_{i,i+1}+(n-1)\frac{J}{2}\mathbb{I},$ (17)
using $\mathrm{SWAP}_{i,j}=(\mathbb{I}+X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j})/2$, and thereby reducing the number of terms in the Hamiltonian by a factor of three.
In general, finding appropriate decompositions is a hard task and thus not necessarily feasible, albeit a trick to be aware of when reducing circuit evaluations. Additionally, non-Hermitian operators, such as the Clifford operators, can not be directly measured and require additional circuitry such as the Hadamard test to perform measurements. As such, using different decompositions might be even more useful in the purely classical simulation regime.
IV Applications
Since the method proposed in Section II.2 relies on preparing an interesting quantum state and applying a function such as time evolution in classical post-processing, we view it as a hybrid quantum-classical algorithm showcasing the power of quantum measurements. Furthermore, any quantum computation will require measurements to output its result in a classically understandable language. Thus, we will always have access to at least a few measurements that could enhance the quantum computation in classical post-processing or as a standalone algorithm where we crank up the number of measurements to improve the accuracy of these methods.
@@ IV.1 Short-time dynamics by measurement
The first practical application lies in quantum simulation. As a guiding example, we have already studied the use for estimating time dynamics of observables via measurements, reminiscent of classical cluster expansion techniques applied to arbitrary quantum states [36]. As such, any application of Cluster expansion, such as the Loschmidt echo $\mathrm{tr}\!\left(\rho\mathrm{e}^{-\mathrm{i}Ht}\right)$, is feasible. Overcoming the restriction to product states and moving towards a hybrid quantum-classical algorithm opens the door to new use cases. For near-term quantum devices, where time evolution algorithms are limited due to device noise but additional measurements might be realistic, we can imagine combining time evolution on the quantum device with additional short-time evolution in classical post-processing to enhance the quantum device capabilities.
To be more concrete, we can imagine a quantum circuit preparing the state $\rho(t_{1})=\mathrm{e}^{-\mathrm{i}Ht_{1}}\rho_{0}\mathrm{e}^{\mathrm{i}Ht_{1}}$ for the Néel state $\rho_{0}=\ket{0,1,0,1,\cdots}\!\bra{0,1,0,1,\cdots}$ | as initial quantum state and the Heisenberg Hamiltonian from Eq. (15). We can then estimate the expectation value of the particle imbalance $O=\sum_{i}(-1)^{i}Z_{i}$ not only for $\rho(t_{1})$ but also classically construct an approximation for $\mathrm{tr}\!\left(O\rho(t_{1}+t_{2})\right)$ for $t_{2}>0$ using the presented method.
@@ IV.2 Short-time Hamiltonian learning and device verification
Another example of practical short-time quantum simulating in post-processing is the task of device verification or Hamiltonian learning [41, 42, 43, 44]. Given a quantum device with an underlying, unknown Hamiltonian, we envision a protocol in which a well-known initial state is prepared and the system then evolves under its Hamiltonian for a short time before the state is measured. In classical post-processing, we can then evolve the system in negative time under a model Hamiltonian and check the statistics on certain observables to test the accuracy of the model Hamiltonian.
@@
We can then gauge the accuracy of a guess $H_{\mathrm{guess}}$ of the system’s Hamiltonian $H_{\mathrm{sys}}$ with the presented method using $\mathrm{tr}\!\left(Z(-t)\rho(t)\right)$ for $Z(-t)=\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}Z\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}$, which would return zero if $H_{\mathrm{sys}}=H_{\mathrm{guess}}$. Repeating this procedure for different input states and observables then provides information about the accuracy of $H_{\mathrm{guess}}$
@@ IV.3 Short-time imaginary time evolution
Apart from real-time evolution, the nested commutator relations also hold for imaginary time evolution, as does the requirement on the truncation order (see, e.g., Ref. [48, Lemma 3.2]). Even though imaginary time evolution is mostly interesting for long simulation times due to its ability to drive a state toward its ground state, even a short simulation can be useful. In a setting where the initial state is prepared in a black-box fashion, e.g., a variational quantum algorithm or the output of some applied shortcut to adiabaticity, even short-time imaginary time evolution will push the prepared state closer toward the system’s ground state. As such, we can, with additional measurements, obtain measurement statistics that are closer to that of the true ground state in classical post-processing.
@@ IV.4 Enhancing sub-universal quantum simulators
The underlying principle of this method is that both the generators of the time evolution operator and the observable belong to the same group. The limitation to short simulation times further means that we are restricted to simulating operators close to the identity. However, in contexts where one cannot perform arbitrary quantum operations, such as sub-universal quantum simulators, even operations close to the identity might be of interest, if they cannot be implemented natively. Therefore, as long as we can measure the generators of a unitary close to the identity and those generators belong to the same group as the observable of interest, we can apply this unitary in classical post-processing and thus expand the capabilities of the quantum simulator.
For example, imagine a quantum device that can only perform gates on neighboring qubits, but measure arbitrary single-qubit Pauli operators. Then, we can add long-range interactions in classical post-processing by, e.g., employing the presented method for $F(t)=\mathrm{e}^{\mathrm{i}H_{1}t_{1}}\cdots\mathrm{e}^{\mathrm{i}H_{r}t_{r}}O\mathrm{e}^{-\mathrm{i}H_{r}t_{r}}\cdots\mathrm{e}^{-\mathrm{i}H_{1}t_{1}}$, where the $H_{i}$ represent shallow sums of geometrically non-local Hermitian operators.
@@
V Discussion
The introduced Taylor series method to estimate short-time dynamics via measurements constitutes another feasible reading of the “measure now, compute functions of interest later” mindset, which combines quantum computations with classical post-processing of measurement statistics. Classical shadows once more present themselves as a multi-faceted tool that can help estimate several unrelated properties simultaneously enabling the search for new post-processing techniques, in this case allowing for a short-time quantum simulation through measurements application.
Although limited to short simulation times, it provides further arguments for the power of quantum measurements, adding to recent insights on the power of Pauli shadows on initial, untrained states in the context of quantum neural networks [49]. In this way, it is in line with other recent developments that aim at fleshing out the power of quantum measurement in quantum state engineering, quantum computation, and quantum simulation. It motivates the further development of hybrid quantum-classical algorithms and new ways to enhance them in post-processing. The limitation to short times is not unexpected since it is present in the purely classical regime of cluster expansion techniques and due to the classical hardness of time evolution in general.
The focus on Pauli operators is not required and decompositions of Hamiltonians into other operators following a group structure, such as Clifford operators, provides another path towards reducing resource costs and simplifying post-processing tasks. We hope the present work stimulates further rigorous research on quantum dynamics and hybrid quantum-classical algorithms.
Acknowledgments
@@ Appendix A Tail bound for the truncated Taylor series
This appendix provides details of arguments presented in the main text. Using insights from the “Trotter error with 1-norm scaling” lemma of Ref. [39, Lemma 1], we can bound the operator norm of the remainder of the truncated Taylor series of the time-evolution operator by using Taylor’s remainder theorem with
@@
recovering the standard bound from Ref. [27, Eq. (4)]. When using $r$ segments of time evolution operators $U(t/r)$ with $r\geq\lVert H\rVert t$, the accuracy required for each segment is $\epsilon/r$, resulting in
@@ Appendix B Direct Taylor expansion of $O(t)$
In Section II.2, we have constructed an approximation of $O(t)$ by concatenating two truncated Taylor series of the time evolution operator. Consequently, the resulting approximation will also contain terms with higher powers of $t$ than the cutoff $K$. Although such an approximation will be more accurate, it differs from a direct Taylor expansion of the target function $F(H,t)=O(t)$. We will now showcase such a direct approach for the example of time dynamics and the resulting trade-off in accuracy and the number of terms in the linear combination of Pauli operators setting $r=1$ for simplicity.
@@ B.1 Expansion
The number of terms in this linear combination of Pauli operators is upper-bounded by $\sum_{k+k^{\prime}\leq K}L_{H}^{k+k^{\prime}}=\mathcal{O}(L^{K})$ and thus quadratically smaller than when concatenating two Taylor approximations. As such, we cannot expect the same accuracy of this approximation and must construct an appropriate bound on the systematic approximation error.
@@ B.2 Error bound
We here present a rigorous error bound. To this end, we can use the “Trotter error with 1-norm scaling” lemma of Ref. [39]. This theorem bounds the tail $\mathrm{\textbf{R}}_{K}$ of product formulas above $K$-th order via
@@
where we have lifted $O$ to fit the product formula form and recovered $\lVert O\rVert$ since Ref. [39, Lemma 6] has an additional factor of $\lVert\prod_{\ell=1}^{L_{H}}\mathrm{e}^{\mathrm{i}\alpha_{\ell}H_{\ell}t}\rVert$ which vanishes for all anti-Unitary operators in the exponent but remains for $\mathrm{e}^{\ln O}$. The error is thus larger by a factor $32^{k+1}/$ 3 compared to the concatenated approach discussed in Section II.3.1. Again, these results could be extended to a commutator bound using insights from Ref. [39].

### v2

@@ III Concrete example and reducing terms
To gain better intuition of how the expansion of time dynamics into a linear combination of Pauli operators of Eq. (5) might look for a specific Hamiltonian, let us consider a Heisenberg Hamiltonian with equally weighted interactions, i.e.,
@@
Naively, we have upper-bounded the number of terms present in the expansion of the $K$-th order truncated Taylor series of the time evolution operator as $m_{\mathrm{tot}}=\sum_{k=0}^{K}L^{k}=(L^{K+1}-1)/(L-1)$. Also, since $\lambda=3nJ$, we have $\lVert\gamma\rVert_{1}=\mathrm{poly}(\mathrm{e}^{(n-1)Jt})$, explicitly showing that the system size has a dependence on the number of required samples and practically feasible simulation time. However, since the Pauli operators square to the identity and form a group, many terms will collapse to the identity matrix or reproduce existing terms as early as in the second order. Explicitly, for
@@
Thus, using algebraic properties of the Pauli operators ensures that the total number of contributing terms for a second-order Taylor series is smaller than the worst-case bound of $(27(n-1)^{3}-1)/(3(n-1)-1)$ by $(n+9)(n-1)$, leading to $m_{\mathrm{tot}}=8n^{2}-23n+16$ terms. Note, however, that although these terms do not contribute to the number of Pauli operators, they still contribute their pre-factors. These pre-factors $\gamma_{j}$ in Eq. (5) will then be sums of powers of $Jt$ and must be considered in the error analysis, ultimately contributing to the sample complexity, where contributions to the identity term and negative contributions reduce $\lVert\gamma\rVert_{1}$ and thus allow for larger simulation times.
However, this constitutes only a modest asymptotic reduction that cannot be explicitly generalized.
@@
$H=\sum_{i=1}^{n-1}\frac{J}{2}\mathrm{SWAP}_{i,i+1}-(n-1)\frac{J}{2}\mathbb{I},$ (19)
using $\mathrm{SWAP}_{i,j}=(\mathbb{I}+X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j})/2$, thereby reducing the number of terms in the Hamiltonian by a factor of three. Finding appropriate approximations is thus essential in improving the practicality of this method.
In general, finding such decompositions is a hard task and thus not necessarily feasible, albeit a trick to be aware of when reducing circuit evaluations. Additionally, non-Hermitian operators, such as the Clifford operators, can not be directly measured and require additional circuitry, such as the Hadamard test, to perform measurements. As such, using different decompositions might be even more useful in the purely classical simulation regime.
IV Discussion of potential applications
The proposed method offers a tradeoff between a genuine quantum operation, such as time evolution, and a more involved classical post-processing stage with additional sample complexity. This can be viewed as a hybrid quantum-classical algorithm that builds on the power of quantum measurements. Furthermore, any quantum computation will require measurements to output its result in a classically understandable language. Thus, we will always have access to at least a few measurements that could enhance most quantum computations in classical post-processing. Alternatively, when implemented as a standalone algorithm, we can crank up the number of measurements to improve its error. However, we think that this method is best applied in combination with a capable but limited quantum device. There, it could be integrated by exhausting the capabilities of the device and enhancing the computational power by additional measurements and the application of the Taylor expansion of a suitable spectral function in post-processing.
@@ IV.1 Short-time dynamics by measurement
The first potential application lies in quantum simulation. As a guiding example, we have already studied the use for estimating time dynamics of observables via measurements, reminiscent of classical cluster expansion techniques applied to arbitrary quantum states [41]. As such, any application of Cluster expansion, such as the Loschmidt echo $\mathrm{tr}\!\left(\rho\mathrm{e}^{-\mathrm{i}Ht}\right)$, is feasible. Overcoming the restriction to product states and moving toward a hybrid quantum-classical algorithm opens the door to new use cases. For near-term quantum devices, where time evolution algorithms are limited due to device noise but additional measurements might be realistic, we can imagine combining time evolution on the quantum device with additional short-time evolution in classical post-processing to enhance the quantum device capabilities.
To be more concrete, we can imagine a quantum circuit preparing the state $\rho(t_{1})=\mathrm{e}^{-\mathrm{i}Ht_{1}}\rho_{0}\mathrm{e}^{\mathrm{i}Ht_{1}}$ for the Néel state $\rho_{0}=\ket{0,1,0,1,\cdots}\!\bra{0,1,0,1,\cdots}$ | as initial quantum state and the Heisenberg Hamiltonian from Eq. (17).
We can then estimate the expectation value of the particle imbalance $O=\sum_{i}(-1)^{i}Z_{i}$ not only for $\rho(t_{1})$ but also classically construct an approximation for $\mathrm{tr}\!\left(O\rho(t_{1}+t_{2})\right)$ for $t_{2}>0$ using the presented method.
As such, we envision a Schrödinger evolution to be implemented on actual quantum hardware and a short-time Heisenberg evolution simulated classically via additional measurements.
Again, we stress that short times really mean short times since the sampling overhead grows as $\mathrm{poly}(\mathrm{e}^{\lambda t})$.
@@ IV.2 Short-time Hamiltonian learning and device verification
Another example of practical short-time quantum simulating in post-processing is the task of device verification or Hamiltonian learning [53, 54, 55, 56]. Given a quantum device with an underlying, unknown Hamiltonian, we envision a protocol in which a well-known initial state is prepared, and the system then evolves under its Hamiltonian for a short time before the state is measured. In classical post-processing, we can then evolve the system in negative time under a model Hamiltonian and check the statistics on certain observables to test the error of the model Hamiltonian.
@@
We can then gauge the error of a guess $H_{\mathrm{guess}}$ of the system’s Hamiltonian $H_{\mathrm{sys}}$ with the presented method using $\mathrm{tr}\!\left(Z(-t)\rho(t)\right)$ for $Z(-t)=\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}Z\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}$, which would return zero if $H_{\mathrm{sys}}=H_{\mathrm{guess}}$. Repeating this procedure for different input states and observables then provides information about the error of $H_{\mathrm{guess}}$
@@ IV.3 Short-time imaginary time evolution
Apart from real-time evolution, the nested commutator relations also hold for imaginary time evolution, as does the requirement on the truncation order (see, e.g., Ref. [60, Lemma 3.2]). Even though imaginary time evolution is mostly interesting for long simulation times due to its ability to drive a state toward its ground state, even a short simulation can be useful. In a setting where the initial state is prepared in a black-box fashion, e.g., a variational quantum algorithm or the output of some applied shortcut to adiabaticity, even short-time imaginary time evolution will push the prepared state closer toward the system’s ground state. As such, we can, with additional measurements, obtain measurement statistics that are closer to that of the true ground state in classical post-processing.
@@
Further, since the 1-norm of the pre-factors of the expansion $\lVert\gamma\rVert_{1}$ is unaffected by moving from real to imaginary time, the sampling overheads are identical for both cases.
@@ IV.4 Enhancing sub-universal quantum simulators
The underlying principle of this method is that both the generators of the time evolution operator and the observable belong to the same group. The limitation to short simulation times further means that we are restricted to simulating operators close to the identity. However, in contexts where one cannot perform arbitrary quantum operations, such as sub-universal quantum simulators, even operations close to the identity might be of interest if they cannot be implemented natively. Therefore, as long as we can measure the generators of a unitary close to the identity and those generators belong to the same group as the observable of interest, we can apply this unitary in classical post-processing and thus expand the capabilities of the quantum simulator.
For example, imagine a quantum device that can only perform gates on neighboring qubits but measure arbitrary single-qubit Pauli operators. Then, we can add long-range interactions in classical post-processing by, e.g., employing the presented method for $F(t)=\mathrm{e}^{\mathrm{i}H_{1}t_{1}}\cdots\mathrm{e}^{\mathrm{i}H_{r}t_{r}}O\mathrm{e}^{-\mathrm{i}H_{r}t_{r}}\cdots\mathrm{e}^{-\mathrm{i}H_{1}t_{1}}$, where the $H_{i}$ represent shallow sums of geometrically non-local Hermitian operators.
@@
V Conclusion and outlook
The introduced Taylor series method to estimate short-time dynamics via measurements constitutes another feasible reading of the “measure now, compute functions of interest later” mindset, which combines quantum computations with classical post-processing of measurement statistics. Classical shadows once more present themselves as a multi-faceted tool that can help estimate several unrelated properties simultaneously, enabling the search for new post-processing techniques, in this case allowing for a short-time quantum simulation through measurement application. We view this method as an enhancement of quantum processors limited in their computational power, e.g., by combining a classically difficult computation such as time evolution where the maximal evolution time is limited by the quantum device with an additional short-time evolution through additional measurements.
Although limited to short simulation times, it provides further arguments for the power of quantum measurements, adding to recent insights on the power of Pauli shadows on initial, untrained states in the context of quantum neural networks [61]. This is in line with other recent developments that aim to flesh out the power of quantum measurement in quantum state engineering, quantum computation, and quantum simulation. It motivates the further development of hybrid quantum-classical algorithms and new ways to enhance them in post-processing. The limitation to short times is not unexpected since it is present in the purely classical regime of cluster expansion techniques and due to the classical hardness of time evolution in general.
The focus on Pauli operators is not required, and decompositions of Hamiltonians into other operators following a group structure, such as Clifford operators, provides another path toward reducing resource costs and simplifying post-processing tasks. We hope the present work stimulates further rigorous research on quantum dynamics and hybrid quantum-classical algorithms.
Note added:
During the review process of this manuscript, we learned of a complementary framework by Fuller et al. [62], which appeared after our first preprint and discusses the integration of classical simulation of backpropagated Heisenberg evolution with Schrödinger evolution executed on quantum hardware. While a detailed comparison of both approaches is of great practical interest, it is outside the scope of this manuscript.
Acknowledgments:
@@ Appendix A Tail bound for the truncated Taylor series
This appendix provides details of arguments presented in the main text. Using insights from the “Trotter error with 1-norm scaling” lemma of Ref. [50, Lemma 1], we can bound the operator norm of the remainder of the truncated Taylor series of the time-evolution operator by using Taylor’s remainder theorem with
@@
recovering the standard bound from Ref. [32, Eq. (4)]. When using $r$ segments of time evolution operators $U(t/r)$ with $r\geq\lVert H\rVert t$, the error required for each segment must be smaller than $\epsilon/r$, resulting in
@@ Appendix B Direct Taylor expansion of $O(t)$
In Section II.2, we have constructed an approximation of $O(t)$ by concatenating two truncated Taylor series of the time evolution operator. Consequently, the resulting approximation will also contain terms with higher powers of $t$ than the cutoff $K$. Although such an approximation will be more accurate, it differs from a direct Taylor expansion of the target function $F(H,t)=O(t)$. We will now showcase such a direct approach for the example of time dynamics and the resulting trade-off in error and the number of terms in the linear combination of Pauli operators setting $r=1$ for simplicity.
@@ B.1 Expansion
The number of terms in this linear combination of Pauli operators is upper-bounded by $\sum_{k+k^{\prime}\leq K}L_{H}^{k+k^{\prime}}=\mathcal{O}(L^{K})$ and is thus quadratically smaller than when concatenating two Taylor approximations. As such, we cannot expect the same error of this approximation and must construct an appropriate bound on the systematic approximation error.
@@ B.2 Error bound
We present a rigorous error bound. To this end, we can use the “Trotter error with 1-norm scaling” lemma of Ref. [50]. This theorem bounds the tail $\mathrm{\textbf{R}}_{K}$ of product formulas above $K$-th order via
@@
where we have lifted $O$ to fit the product formula form and recovered $\lVert O\rVert$ since Ref. [50, Lemma 6] has an additional factor of $\lVert\prod_{\ell=1}^{L_{H}}\mathrm{e}^{\mathrm{i}\alpha_{\ell}H_{\ell}t}\rVert$ which vanishes for all anti-Unitary operators in the exponent but remains for $\mathrm{e}^{\ln O}$. The error is thus larger by a factor $32^{k+1}/$ 3 compared to the concatenated approach discussed in Section II.3.1. Again, these results could be extended to a commutator bound using insights from Ref. [50].